### 30 Apr 2026 v2

# Auditing Marketing Budget Allocation with Hindsight Regret

Abstract.
Organizations routinely make strategic budget allocations under operational constraints. However, they often lack a principled way to assess those realized allocations. We present a retrospective auditing framework based on hindsight regret, for marketing budget allocation. Regret is defined as the opportunity cost of the realized allocation relative to a constraint-faithful benchmark under the same budget and stability guardrails. The framework estimates regime-specific spend–response functions from historical logs, computes feasible hindsight allocations by constrained optimization, and propagates uncertainty through Monte Carlo evaluation to produce regret distributions, expected lift, and probability-of-improvement summaries. This separates allocation inefficiency from uncertainty in the estimated response surfaces. Experiments on real marketing allocation logs show that the framework yields interpretable post-hoc diagnostics and reveals a practical trade-off between allocation flexibility and detectability: moderate feasible re-allocations often capture most measurable gain, while larger shifts move into weak-support regions with higher uncertainty. The result is a practical method for auditing historical budget decisions when online experimentation is costly or infeasible.
Regret Analysis, Decision Theory, Marketing Portfolio Allocation, Causal Inference , Counterfactual Evaluation, Partial Observability
$c$s: Computing methodologies Causal reasoning and diagnosticsccs: Computing methodologies Machine learningccs: Information systems Online advertisingccs: Computing methodologies Uncertainty quantificationccs: Computing methodologies Continuous optimization
1. Introduction
Strategic budget allocation is a recurring decision problem in digital advertising. Organizations must distribute finite budget across channels, markets, and campaigns over a planning horizon while respecting operational guardrails such as budget limits, pacing rules, and stability constraints. Because these allocations are high-dimensional, persistent, and costly to revise, poor decisions can carry substantial economic consequences.
A central challenge is retrospective evaluation. Observed outcomes such as revenue or profit reflect not only decision quality, but also market conditions, platform dynamics, and other factors outside the planner’s control. As a result, even when historical logs are available, it is difficult to determine whether a realized allocation was close to the best feasible decision that could have been made under the same operational constraints. This raises a natural question: How much value was left on the table due to the allocation decision itself?
Online advertising provides a concrete and economically important instance of this problem. Although auctions operate at the level of individual impressions, many consequential choices occur at a slower timescale through portfolio-style budget commitments across heterogeneous assets such as channels, regions, and campaigns (Zhao et al., 2019). These decisions are subject to practical constraints including caps, pacing rules, and limits on reallocation across periods (Balseiro et al., 2024; Nguyen et al., 2023; Holthausen and Assmus, 1982; Symitsi et al., 2022; Baardman et al., 2019). In practice, organizations often have detailed historical allocation logs, but lack principled tools for auditing whether past allocations were near-optimal.
Standard evaluation approaches do not directly address this need. Randomized experiments are costly or infeasible for continuous, high-dimensional allocations and may violate no-interference assumptions due to cross-channel effects (Kohavi et al., 2020; Eckles et al., 2017). Pre–post and interrupted time-series analyses are sensitive to non-stationarity and shocks (Bernal et al., 2017; Angrist and Pischke, 2009). Synthetic control methods rely on stable donor units and cross-sectional comparability assumptions that are difficult to justify for persistent, multi-asset allocation trajectories (Abadie, 2021; Dubé and Zipperer, 2014). More broadly, while regret is central in online learning, those formulations focus on sequential adaptation under partial feedback rather than post-hoc auditing of a realized strategic allocation.
We therefore study retrospective budget-allocation auditing through hindsight regret: the opportunity cost of the realized allocation relative to the best feasible alternative in hindsight under the same operational guardrails. Our goal is not online control or policy learning, but post-hoc measurement of whether a historical allocation left recoverable value unrealized. To do so, we estimate regime-specific spend–response functions from historical logs, construct constraint-faithful hindsight benchmarks through constrained optimization, and propagate model and outcome uncertainty through Monte Carlo evaluation to obtain regret distributions and probability-of-improvement summaries. This yields an interpretable estimate of “value left on the table” while separating apparent allocation inefficiency from uncertainty in the estimated response surfaces.
In summary, our main contributions are:
  1. (1)
    A hindsight-regret formulation for retrospective allocation auditing. We cast post-hoc evaluation of strategic budget decisions as comparison against a constraint-faithful hindsight benchmark, yielding an interpretable measure of opportunity cost.
  2. (2)
    A practical offline auditing framework. We combine regime-specific spend–response estimation, constrained hindsight optimization, and uncertainty propagation to evaluate historical allocations using logged data alone.
  3. (3)
    An empirical study on real strategic allocation logs. We show how regret distributions, detectability thresholds, and stability constraints together support systematic post-hoc assessment of budget decisions.
2. Background & Related Work
Overview of Measurement Approaches in Practice: Randomized online experiments remain the standard for causal measurement in digital systems (Kohavi et al., 2020). However, strategic budget allocation is continuous, high-dimensional, and persistent, which makes large-scale randomization costly and often operationally infeasible. Advertising systems also exhibit interference across auctions, channels, and users, weakening standard no-interference assumptions (Eckles et al., 2017). When experimentation is unavailable, practitioners often fall back on pre–post or interrupted time-series designs (Bernal et al., 2017; Angrist and Pischke, 2009), which rely on stable untreated dynamics and can break under demand shocks or non-stationarity. Synthetic control methods similarly require donor stability and cross-sectional comparability (Abadie, 2021; Dubé and Zipperer, 2014; Zhang, 2019), assumptions that are difficult to justify for persistent, policy-driven allocation trajectories in mediated ad-delivery systems.
Regret as retrospective decision evaluation: In statistical decision theory, regret formalizes an ex-post opportunity cost: given realized outcomes and a feasible action set, the chosen decision is evaluated against the best feasible alternative that could have been selected in hindsight under the same environment (Savage, 1954; Loomes and Sugden, 1982). This produces a static optimality gap that measures decision inefficiency rather than learning dynamics. By contrast, regret in bandit and reinforcement learning settings (Auer, 2002; Jaksch et al., 2010; Srinivas et al., 2010; Abbasi-Yadkori et al., 2011; Dean et al., 2018) is sequential and cumulative, reflecting the loss incurred while adaptively learning under partial feedback. Our formulation adopts the former perspective. It is also distinct from off-policy evaluation, whose objective is to estimate the value of a target policy from logged feedback. Here the object of evaluation is a realized allocation trajectory, and the benchmark is a feasible hindsight reallocation under the same operational regime rather than a new policy to be deployed prospectively.
Structural Causal Modeling Approach: Budget allocation in advertising systems operates through a structural data-generating process: planned budgets affect latent bidding and delivery decisions, which in turn determine realized spend and downstream returns (Bottou et al., 2013). We use the Structural Causal Model (SCM) framework (Pearl, 2009) to give semantics to these counterfactual comparisons through interventional response functions, namely expected returns under hypothetical spend allocations. Identifiability of such structural response functions requires assumptions on functional form and noise structure. Prior work on additive and related noise models provides identifiability results for nonlinear cause–effect pairs (Shimizu et al., 2006; Hoyer et al., 2009; Peters et al., 2017; Immer et al., 2023; Yin et al., 2024).
Need for a Custom Framework: Existing approaches do not directly answer the question central to this paper: whether a realized strategic allocation was close to the best feasible allocation that could have been chosen in hindsight under the same operational guardrails. Our framework addresses this gap by estimating regime-specific response surfaces from historical logs, constructing feasible hindsight reallocations within the same regime, and quantifying the resulting opportunity cost with uncertainty-aware regret summaries. Non-stationarity is handled through epoch-wise re-estimation and support-aware uncertainty inflation rather than by assuming invariant untreated baselines or stationary learning dynamics.
3. Problem Formulation
We consider sequential budget allocation over a fixed planning horizon of $T$ operational time steps. Budgets are updated at coarser decision epochs $\tau\in\{1,\dots,T/k\}$ and held fixed for $k$ operational steps. At each epoch, the planner allocates budget across $N$ assets.
Let $\mathbf{b}_{\tau}=(b_{\tau,1},\dots,b_{\tau,N})\in\mathcal{B}\subseteq\mathbb{R}_{+}^{N}$ denote the epoch-$\tau$ budget vector, where $\mathcal{B}$ encodes feasibility constraints such as total budget, caps, floors, and operational guardrails.
Mediated execution.
Budgets do not directly determine spend or returns. Instead, delivery is mediated by a platform execution mechanism interacting with a stochastic market environment (Bottou et al., 2013). Let $x_{\tau}$ denote observed context, $u_{\tau}$ unobserved exogenous factors (e.g., competition, inventory), and $z_{\tau}$ internal system state. The structural relations are
(1) $\boldsymbol{\beta}_{\tau}$ $=\pi(\mathbf{b}_{\tau},x_{\tau},z_{\tau},u_{\tau}),$
(2) $\mathbf{s}_{\tau}$ $=\sigma(\boldsymbol{\beta}_{\tau},x_{\tau},z_{\tau},u_{\tau})+\boldsymbol{\xi}_{\tau},\quad\mathbb{E}[\boldsymbol{\xi}_{\tau}|\cdot]=0,$
(3) $\boldsymbol{\rho}_{\tau}$ $=\rho(\mathbf{s}_{\tau},x_{\tau},u_{\tau})+\boldsymbol{\varepsilon}_{\tau},\quad\mathbb{E}[\boldsymbol{\varepsilon}_{\tau}|\cdot]=0.$
Here $\mathbf{s}_{\tau}$ denotes realized spend and $\boldsymbol{\rho}_{\tau}$ realized returns. Latent bidding and serving actions $\boldsymbol{\beta}_{\tau}$ are not observed. Our retrospective analysis therefore operates at the realized-spend layer, treating the upstream execution mechanism as latent.
Observability and feedback.
At epoch $\tau$ the planner observes $(x_{\tau},\mathbf{b}_{\tau},\mathbf{s}_{\tau},\boldsymbol{\rho}_{\tau})$, but not $\boldsymbol{\beta}_{\tau}$, $z_{\tau}$, $u_{\tau}$, or counterfactual outcomes. Feedback is bandit-style, revealing returns only at the executed allocation.
Budget accounting.
The planner is given a fixed total budget $B_{\mathrm{tot}}$ for the horizon, with allocations satisfying
(4) $\sum_{\tau=1}^{T/k}\sum_{i=1}^{N}b_{\tau,i}=B_{\mathrm{tot}},\qquad\mathbf{b}_{\tau}\in\mathcal{B}\ \forall\tau.$
Planning policy.
A policy $\Pi$ maps available information to allocations: $\mathbf{b}_{\tau}\leftarrow\Pi(\mathcal{F}_{\tau})$, with $\mathcal{F}_{\tau}=\{x_{1:\tau},\mathbf{b}_{1:\tau-1},\mathbf{s}_{1:\tau-1},\boldsymbol{\rho}_{1:\tau-1}\}.$ This partially observed sequential formulation defines the historical setting against which hindsight regret is evaluated.
4. Retrospective Regret Framework
We present a retrospective framework for auditing the quality of a realized budget allocation through hindsight regret, defined as the opportunity cost relative to the best feasible spend trajectory that could have been executed under the same operational constraints and historical regime. The pipeline proceeds in three steps:
  1. (1)
    Interventional spend–return modeling (Step I): Estimate uncertainty aware interventional spend–return response functions under a reduced-form grey-box abstraction.
  2. (2)
    Constraint-faithful oracle allocation (Step II): Compute the best feasible benchmark spend trajectory under operational guardrails via constrained optimization.
  3. (3)
    Uncertainty-aware regret estimation (Step III): Propagate response uncertainty by Monte Carlo simulation to obtain a distribution over regret and associated summary metrics.
4.1. Empirical regret
In classical decision theory, regret is the loss in utility incurred by a decision that is suboptimal relative to an alternative evaluated in hindsight under the realized state of the world (Savage, 1954; Loomes and Sugden, 1982). In our setting, a decision corresponds to a feasible spend trajectory $\mathbf{s}=\{s_{\tau,i}\}_{\tau,i}$ and utility is total realized return $R(\mathbf{s}):=\sum_{\tau,i}\rho_{\tau,i}(s_{\tau,i})$. Given a realized trajectory $\mathbf{s}_{\mathrm{real}}$ and a feasible benchmark $\mathbf{s}^{\star}$, regret is defined as
(5) $\mathrm{Reg}:=R(\mathbf{s}^{\star})-R(\mathbf{s}_{\mathrm{real}}).$
Because realized returns are stochastic and response functions are estimated from data, $\mathrm{Reg}$ is a random variable. Individual realizations may be negative, corresponding to favorable stochastic outcomes under the realized trajectory relative to the benchmark. We therefore evaluate expected regret
(6) $\mathbb{E}[\mathrm{Reg}]=\mathbb{E}\!\left[R(\mathbf{s}^{\star})-R(\mathbf{s}_{\mathrm{real}})\right],$
and report probability statements of the form $\mathbb{P}(\mathrm{Reg}>0)$, which quantify the likelihood that the benchmark dominates the realized trajectory in hindsight.
Operational regret with stability constraints ($\delta$) and confidence ($\varepsilon$).
The benchmark depends on the operational realism encoded by the feasible set. We index the feasible set by a stability parameter $\delta$ and define
(7) $\mathrm{Reg}(\delta):=R(\mathbf{s}^{\star}(\delta))-R(\mathbf{s}_{\mathrm{real}}),\qquad\mathbf{s}^{\star}(\delta)\in\arg\max_{\mathbf{s}\in\mathcal{B}(\delta)}R(\mathbf{s}),$
where $\mathcal{B}(\delta)$ denotes the $\delta$-constrained feasible set (defined in Step II). We call regret $\varepsilon$-certified at level $\delta$ if
(8) $\mathbb{P}\!\left(\mathrm{Reg}(\delta)>0\right)\geq\varepsilon.$
Let $\Delta$ denote a prespecified set of plausible constraint levels. We select
(9) $\delta^{\star}\in\arg\max_{\delta\in\Delta}\mathbb{E}\!\left[\mathrm{Reg}(\delta)\right]\quad\text{s.t.}\quad\mathbb{P}\!\left(\mathrm{Reg}(\delta)>0\right)\geq\varepsilon,$
and report $\mathbb{E}[\mathrm{Reg}(\delta^{\star})]$ together with $\mathbb{P}(\mathrm{Reg}(\delta^{\star})>0)$.
Metric Definition Interpretation
Expected regret $\widehat{\mathbb{E}}[\mathrm{Reg}(\delta^{\star})]=\frac{1}{J}\sum_{j=1}^{J}\mathrm{Reg}^{(j)}(\delta^{\star})$ Average achievable gain from feasible hindsight reallocation at the selected constraint level.
Standard deviation $\widehat{\sigma}_{\mathrm{Reg}}=\sqrt{\frac{1}{J-1}\sum_{j}(\mathrm{Reg}^{(j)}(\delta^{\star})-\widehat{\mathbb{E}}[\mathrm{Reg}(\delta^{\star})])^{2}}$ Sensitivity of regret to outcome noise and model uncertainty; large values indicate fragile conclusions.
Credible interval $\mathrm{CI}_{1-\alpha}=[Q_{\alpha/2},Q_{1-\alpha/2}]$ Range of plausible regret values under uncertainty; intervals spanning zero indicate ambiguous dominance.
Probability of improvement $\widehat{\mathbb{P}}(\mathrm{Reg}(\delta^{\star})>0)=\frac{1}{J}\sum_{j}\mathds{1}\{\mathrm{Reg}^{(j)}(\delta^{\star})>0\}$ Likelihood that the $\delta^{\star}$-constrained benchmark outperforms the realized trajectory; compared against $\varepsilon$ in (8).
Table 1. Empirical regret metrics and uncertainty summaries. All quantities are computed from Monte Carlo draws $\{\mathrm{Reg}^{(j)}(\delta)\}_{j=1}^{J}$. Unless otherwise stated, we evaluate at $\delta=\delta^{\star}$ selected by (9).
4.2. Step I: Grey-Box Response Functions
We model epoch-specific spend–return mappings and evaluate counterfactual spend trajectories under fixed operational guardrails. Upstream budget-to-bid dynamics are abstracted away; analysis is conducted at the realized-spend layer.
4.2.1. Structural Target and Identification
We formalize the assumptions used to identify an epoch-specific interventional mean response. The key modeling choice is to treat each epoch as a locally stationary regime: within an epoch, observations are analyzed as conditionally independent replicates given observed covariates, while non-stationarity is handled at epoch boundaries through drift regularization and stability constraints.
A1. Structural Response Model.
For asset $i$ in epoch $e$, realized outcomes satisfy
(10) $\rho_{\tau,i}=f_{i,e}(s_{\tau,i},x_{\tau,i})+\varepsilon_{\tau,i},\qquad\tau\in e.$
The function $f_{i,e}$ represents the reduced-form spend–return mechanism induced by the execution regime in epoch $e$.
A2. Within-epoch stationarity and local mean independence (identification).
Within epoch $e$, conditional on $(s_{\tau,i},x_{\tau,i})$, the noise process $\{\varepsilon_{\tau,i}\}_{\tau\in e}$ is treated as mean-zero and approximately independent across $\tau$:
(11) $\mathbb{E}[\varepsilon_{\tau,i}\mid s_{\tau,i},x_{\tau,i}]=0,\qquad\varepsilon_{\tau,i}\ \perp\!\!\!\perp\ \varepsilon_{\tau^{\prime},i}\ \ (\tau\neq\tau^{\prime};\ \tau,\tau^{\prime}\in e).$
Under A2, the interventional mean response
$\mu_{i,e}(s)=\mathbb{E}\!\left[f_{i,e}(s,X_{\tau,i})\mid\tau\in e\right]$
is identified over the realized spend range in epoch $e$, that is, the $s$ values observed under the historical policy within that epoch.
A3. Bounded drift across epochs (regularization only).
Across adjacent epochs $e$ and $e^{\prime}$, the reduced-form mechanism varies smoothly on the overlap of their realized spend ranges:
(12) $\sup_{s\in\mathcal{S}_{e,e^{\prime}}}|f_{i,e}(s)-f_{i,e^{\prime}}(s)|\leq\delta,\qquad\mathcal{S}_{e,e^{\prime}}:=\mathrm{range}(e)\cap\mathrm{range}(e^{\prime}).$
This condition enables weak borrowing of strength for estimation stability but does not expand causal identification beyond the epoch’s realized spend range. Accordingly, the framework is intended for within-regime retrospective auditing, with conservative uncertainty-aware evaluation when optimization reaches weak-support regions.
4.2.2. Drift-regularized grey-box model
The mean response is decomposed as
(13) $\rho_{\tau,i}=\mu_{\mathrm{nom},i,e}(s_{\tau,i};\theta_{i,e})+g_{i,e}(s_{\tau,i})+\sigma_{i,e}(s_{\tau,i})\,\varepsilon_{\tau,i}.$
Nominal saturation.
(14) $\mu_{\mathrm{nom},i,e}(s)=a_{i,e}(1-e^{-b_{i,e}s}),\qquad a_{i,e},b_{i,e}\geq 0.$
Residual mean and variance.
$g_{i,e}(\cdot)\sim\mathcal{GP}(0,k_{g}),\qquad\log\sigma^{2}_{i,e}(\cdot)\sim\mathcal{GP}(m_{\sigma},k_{\sigma}),$
with Matérn(2.5)(2.5) kernels.
The two Gaussian processes serve distinct purposes. The mean GP $g_{i,e}$ captures systematic deviations from the nominal saturation curve, allowing flexible local curvature while preserving the interpretable global shape imposed by $\mu_{\mathrm{nom},i,e}$. The variance GP models heteroskedastic outcome dispersion as a function of spend, enabling uncertainty to increase in weak-support regions. Separating mean and variance components yields calibrated counterfactual uncertainty, which is essential for Monte Carlo regret estimation and detectability analysis.
Auxiliary regularization.
Neighboring-epoch samples are incorporated with exponential weights
$w_{t}=\alpha_{\mathrm{aux}}\exp(-|d_{t}|/\tau),$
which is equivalent to treating auxiliary points as noisier observations. Core epoch data determine identification; auxiliary data provide weak curvature regularization under bounded drift.
Support-aware extrapolation.
For $s$ beyond the upper boundary of the core epoch’s realized spend range, the fitted mean is projected onto the space of non-decreasing functions via boundary-anchored isotonic regression. For $s$ beyond the weighted realized range (core plus down-weighted auxiliary coverage), epistemic variance is inflated:
(15) $\sigma_{\mathrm{epi}}^{2}(s)=\sigma_{g}^{2}(s)+\left[\kappa_{\mathrm{right}}\left(\frac{\max\{0,s-s_{\mathrm{hi}}^{\mathrm{all}}\}}{s_{\mathrm{scale}}}\right)^{p}\sigma_{\mathrm{res}}\right]^{2}.$
These regions are used only for conservative regret evaluation. Within the core realized range, monotonicity is not imposed; in practice, the posterior mean is typically monotone due to the dominance of the nominal saturation component and smooth GP regularization. Example of the constructed oracle functional forms are shown in Figure 1.
4.3. Step II: Oracle allocation under operational feasibility constraints
Figure 1. Monotone grey-box spend–response estimation under bounded temporal drift. Current-period observations define the core realized spend range (blue), while adjacent periods (grey) provide weakly weighted curvature information. Posterior uncertainty widens outside the core range to reflect increased epistemic risk.
Given the fitted epoch-wise response functions $\{\hat{\mu}_{i,e}(\cdot)\}_{i=1}^{K}$, we define an oracle allocation as the solution to a fixed objective evaluated over alternative feasible sets:
(16) $\max_{\{s_{\tau,i}\}_{\tau,i}}\quad\sum_{e=1}^{E}\sum_{i=1}^{K}\hat{\mu}_{i,e}(s_{e,i}),$
where $s_{e,i}$ denotes the spend assigned to asset $i$ in epoch $e$. The distinction between oracle variants arises solely from differences in the feasible set, which encodes alternative assumptions about operational realism.
4.3.1. Unconstrained oracle (idealized hindsight)
The unconstrained oracle evaluates (16) over the largest feasible set $\mathcal{B}_{\mathrm{unc}}$, defined only by budget accounting and asset-level bounds:
$\mathcal{B}_{\mathrm{unc}}:=\left\{\mathbf{s}\;\middle|\;\sum_{e,i}s_{e,i}=B_{\mathrm{tot}},\;\underline{s}_{i}\leq s_{e,i}\leq\overline{s}_{i}\right\}.$
This oracle represents a theoretical upper bound corresponding to perfect hindsight and unrestricted reallocation across epochs.
4.3.2. Constrained oracle (realism-adjusted)
Unconstrained optimization may produce trajectories that are operationally implausible. To restore realism, we introduce a constrained family $\mathcal{B}_{\mathrm{con}}(\delta)\subset\mathcal{B}_{\mathrm{unc}}$ that limits how allocations may change across decision epochs. We impose epoch-to-epoch stability constraints:
(17) $s_{e-1,i}(1-\delta)\leq s_{e,i}\leq s_{e-1,i}(1+\delta),\qquad\forall e>1,\ i,$
where $\delta\in(0,1)$ encodes allowable relative changes.
The constrained oracle is defined as
$\mathbf{s}^{\star}_{\mathrm{con}}(\delta)\in\arg\max_{\mathbf{s}\in\mathcal{B}_{\mathrm{con}}(\delta)}\sum_{e=1}^{E}\sum_{i=1}^{K}\hat{\mu}_{i,e}(s_{e,i}).$
(a) $\delta=30\%$
(b) $\delta=50\%$
(c) Unconstrained
Figure 2. Monte Carlo regret (lift) distributions under alternative stability constraints. Each violin summarizes posterior lift relative to the realized allocation across portfolio–horizon pairs. Increasing flexibility expands the feasible set but does not guarantee monotone improvements in uncertainty-aware lift.
4.3.3. Relationship between oracle variants under support-aware planning
Oracle variants differ only through the feasible set $\mathcal{B}(\delta)$, where larger $\delta$ permits greater epoch-to-epoch reallocation. For the deterministic planning objective $\tilde{R}(\mathbf{s})$ in Step II, nested feasibility implies
$\max_{\mathbf{s}\in\mathcal{B}(\delta_{2})}\tilde{R}(\mathbf{s})\;\geq\;\max_{\mathbf{s}\in\mathcal{B}(\delta_{1})}\tilde{R}(\mathbf{s}),\quad\delta_{2}>\delta_{1}.$
Thus monotonicity holds for the deterministic planning objective, but not necessarily for uncertainty-aware empirical regret after posterior evaluation in Step III.
However, we report uncertainty-aware regret from Step III, where the chosen allocation $\mathbf{s}^{\star}(\delta)$ is evaluated under posterior predictive uncertainty. Larger $\delta$ may shift the optimizer toward weakly supported regions, increasing variance and occasionally reducing both $\mathbb{E}[\mathrm{Reg}(\delta)]$ and the detectability $\mathbb{P}(\mathrm{Reg}(\delta)>0)$. Thus operational lift need not be monotone in $\delta$, and empirical curves may exhibit saturation or dips.
4.4. Step III: Regret distribution via Monte Carlo
Regret is random due to stochastic outcomes and uncertainty in the Step I response surfaces. For each stability level $\delta\in\Delta$ (and the unconstrained case), we first compute the oracle trajectory $\mathbf{s}^{\star}(\delta)$ by optimizing the plug-in mean responses from Step I over the feasible set from Step II. We then hold $\mathbf{s}^{\star}(\delta)$ fixed and propagate uncertainty via Monte Carlo to obtain a distribution over $\mathrm{Reg}(\delta)$. For each scenario we perform the following sampling steps:
  1. (1)
    Sample counterfactual returns. Draw independent innovations $\varepsilon_{e,i}^{(j)}\sim\mathcal{N}(0,1)$ and generate counterfactual returns using the fitted location–scale model,
    $\rho_{e,i}^{(j)}(s)=\hat{\mu}_{i,e}(s)+\eta_{i,e}^{(j)}(s)+\hat{\sigma}_{i,e}(s)\,\varepsilon_{e,i}^{(j)},$
    where $\eta_{i,e}^{(j)}(s)$ is a GP-consistent mean perturbation for $g_{i,e}$, augmented with the off-range inflation term in (15). This simulation treats draws as conditionally independent across assets and epochs given the fitted responses. We treat this as a first-order approximation for retrospective auditing rather than a full joint model of cross-asset dependence.
  2. (2)
    Compute regret at each constraint level. Using the fixed oracle trajectory $\mathbf{s}^{\star}(\delta)$, compute
    $\mathrm{Reg}^{(j)}(\delta)=R^{(j)}\!\left(\mathbf{s}^{\star}(\delta)\right)-R^{(j)}\!\left(\mathbf{s}_{\mathrm{real}}\right),\qquad\forall\delta\in\Delta,$
    and analogously $\mathrm{Reg}^{(j)}_{\mathrm{unc}}$ for the unconstrained oracle.
We then select $\delta^{\star}$ via (9) using these Monte Carlo estimates, including the $\varepsilon$-certificate, and report the regret distribution at $\delta^{\star}$ together with the summary metrics in Table 1.
5. Results
We evaluate the proposed retrospective regret framework using historical strategic allocation logs under operationally realistic feasibility constraints. All outcomes are reported as relative lift with respect to the realized allocation. For confidentiality, we omit portfolio composition, asset identities, absolute magnitudes, and the semantic meaning of evaluation horizons. Our goal is not to recover a ground-truth global optimum, but to assess whether the framework yields stable and decision-relevant retrospective diagnostics under realistic feasibility constraints. We address four questions:
  • RQ1: Do logged allocations support stable and interpretable regret diagnostics?
  • RQ2: How does uncertainty shape regret?
  • RQ3: What is the effect of operational stability constraints?
  • RQ4: When are larger reallocations justified?
Figure 3. Ablation of stability for three portfolios and horizons. Mean lift (dashed line) and lift distributions are shown as a function of allowable inter-epoch change $\delta$. Checkmarks denote $P(\mathrm{Lift}>0)\geq 0.8$. Gains typically concentrate at moderate flexibility and saturate thereafter.
5.1. Retrospective Regret, Uncertainty, and Feasibility
We evaluate $P=8$ portfolios across $H=8$ decision horizons. For each portfolio–horizon pair, epoch-specific spend–return functions are estimated using the weighted monotone grey-box model (Section 4.2). Oracle allocations are computed once per feasibility specification and then held fixed. Uncertainty is introduced only at evaluation time via Monte Carlo propagation of outcome noise and model uncertainty. Figure 2 shows lift distributions for representative portfolios across horizons under multiple feasibility settings. The distributions illustrate that regret is inherently distributional and that its dispersion varies across constraint levels.
RQ1–RQ3 findings. Across all portfolio–horizon pairs we observe:
  1. (1)
    Logged allocations support interpretable regret diagnostics (RQ1). The framework produces well-defined lift distributions and certification rates from historical logs without requiring new experimentation, provided evaluation remains within or near empirically supported regions.
  2. (2)
    Uncertainty is first-order (RQ2). Many cases exhibit wide lift dispersion even when posterior mean lift is positive. Negative realizations arise naturally under Monte Carlo evaluation due to stochastic outcomes and epistemic uncertainty.
  3. (3)
    Operational constraints affect both magnitude and detectability (RQ3). Relaxing the stability constraint $\delta$ typically increases plug-in lift, but can widen regret distributions as optimization moves spend toward regions with weaker empirical support, reducing high-confidence detectability.
Aggregate detectability. Table 2 summarizes certification results across feasibility levels. For each $\delta$ and confidence threshold $\epsilon$, we report the fraction of portfolio–horizon pairs satisfying $P(\mathrm{Reg}>0)\geq\epsilon$ and the mean lift among detectable pairs.
Three consistent patterns emerge. First, moderate thresholds ($\epsilon\in\{0.5,0.6\}$) frequently certify improvement. Second, stringent thresholds ($\epsilon\geq 0.8$) substantially reduce detectability, reflecting broad regret distributions even when mean lift is positive. Third, increasing flexibility generally raises mean lift among detectable cases but does not guarantee monotone gains in detectability, consistent with uncertainty inflation beyond core empirical support. Taken together, these results suggest that the framework is most useful as an auditing tool for detecting recoverable opportunity cost under plausible reallocation limits, rather than for certifying large aggressive reallocations.
Constraint $\delta$ $\epsilon{=}0.6$ $\epsilon{=}0.7$ $\epsilon{=}0.8$ $\epsilon{=}0.9$
$\pm 20\%$ 0.59 (8.2 $\pm$ 5.6) 0.33 (8.3 $\pm$ 5.0) 0.20 (8.2 $\pm$ 2.9) 0.08 (9.6 $\pm$ 1.5)
$\pm 30\%$ 0.75 (9.4 $\pm$ 6.7) 0.53 (10.6 $\pm$ 6.6) 0.28 (10.3 $\pm$ 5.3) 0.09 (11.4 $\pm$ 4.1)
$\pm 40\%$ 0.88 (9.9 $\pm$ 7.6) 0.61 (11.4 $\pm$ 7.6) 0.38 (11.3 $\pm$ 7.0) 0.09 (12.2 $\pm$ 4.4)
$\pm 50\%$ 0.89 (10.5 $\pm$ 8.0) 0.66 (11.9 $\pm$ 8.2) 0.42 (12.8 $\pm$ 8.4) 0.12 (12.7 $\pm$ 4.9)
Unconstrained 0.89 (12.5 $\pm$ 9.7) 0.67 (14.4 $\pm$ 10.0) 0.47 (14.0 $\pm$ 9.5) 0.22 (14.7 $\pm$ 6.8)
Table 2. Fraction of portfolio–horizon pairs with $P(\mathrm{Reg}>0)\geq\epsilon$. Parentheses show mean $\pm$ std lift (%) across detectable pairs.
5.2. When Are Large Re-Allocations Justified?
RQ4 evaluates the marginal benefit of relaxing the inter-epoch stability constraint $\delta$. For each $\delta$, we recompute the oracle allocation and estimate regret under the same Monte Carlo protocol, isolating the effect of flexibility alone.
Figure 3 shows two regimes. Allowing moderate flexibility ($\pm 20\%$$\pm 40\%$) yields large gains by correcting obvious misallocations and shifting spend toward well-supported, high-return regions. Beyond this, lift often saturates: once allocations approach local optima, additional movement mainly explores flatter or weakly supported regions where marginal returns are small and uncertainty is higher.
In some portfolios, lift continues to rise with $\delta$ when profitable regions remain farther from the realized allocation, so larger moves are still beneficial. However, greater flexibility typically increases variance, so detectability improves only selectively and can weaken at stringent confidence thresholds, as reflected in Table 2. Overall, most gains are captured at moderate $\delta$, while extreme reallocations trade reliability for limited additional return.
6. Conclusion and Future Work
We introduced a hindsight-regret framework for retrospective auditing of strategic budget allocation under operational constraints. The framework evaluates realized allocations against constraint-feasible hindsight benchmarks built from regime-specific response surfaces estimated from historical logs. Empirically, most recoverable gains are attained through moderate feasible reallocations, while more aggressive flexibility yields diminishing returns and higher uncertainty.
Future work will develop formal guarantees for response estimation, regret perturbation, and finite-sample Monte Carlo summaries. An important next step is a semi-synthetic benchmark with known response structure and oracle regret, which would enable direct recovery-based validation of the framework. We also plan to extend the same setup to policy-level retrospective scoreboarding of alternative allocation strategies within the same logged regime.